Please answer asap I need this.

A photocopy shop in a monopolistically competitive market could daily sell 5,000 copies at a price of 6 cents, or it could sell 7,500 copies at a price of 5 cents.


a. The marginal revenue associated with this range of the business's demand curve is  cent(s).


b. Because this business is a monopolistic competitor, marginal revenue is  

(Click to select)  price.

   options: less than, equal tom greater than

c. . Draw the relevant range of this business's demand curve and identify a point on its marginal revenue curve (graph).

plesse answer asap before april 20 2019.

 



1
Expert's answer
2021-04-22T07:50:05-0400

(a) demand curve

"\u2206y\/\u2206x"

"\u2206y=5-6=-1"

"\u2206x=7500-5000=2500"

"\frac{-1}{2500}=-0.0004"

"P=b-0.0004Q"

Substitute to get b

"6=b-5000(0.0004)"

"b=8"

Demand curve Will be

"P=8-0.0004Q"


Total Revenue = Price × Quantity

Marginal Revenue = Change in Total Revenue ÷ Change in Quantity

Total Revenue at 5000 Copies = 6 cents × 5000 = 30,000 cents

Total Revenue at 7500 Copies = 5 cents ×7500 = 37,500 cents

Marginal Revenue "= (37,500 - 30,000) \/ (7500 - 5000)"

"=\frac{ 7500 }{ 2500}"

"= 3 cents."

(b) marginal revenue is less than the price.

(C)